Six blue books and three red books in a bag. If three (3) books are selected and replaced. Find the probability that at least one is red.
Answer by stanbon(75887)

Six blue books and three red books in a bag. If three (3) books are selected and replaced. Find the probability that at least one is red.
Prob(at least one red) = 1 - Prob(none red)
Prob( none red) = Prob (blue,blue,blue)= (6/9)^3=0.296...
Therefore Prob (at least one red) = 1-0.296...=0.70...
Cheers,
Stan H.